Which of the following stocks is correctly priced if the risk-free rate of return is 3.2 percent and the market rate of return is 11.76 percent? More than one may be correct

Stock Beta Expected Return
A 0.87 11.03%
B 1.09 12.50%
C 1.18 13.21%
D 1.34 15.02%
E 1.62 17.07%
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Required return for:

A=3.2+(11.76-3.2)*0.87=10.6472%

B=3.2+(11.76-3.2)*1.09=12.5304%

C=3.2+(11.76-3.2)*1.18=13.3008%

D=3.2+(11.76-3.2)*1.34=14.6704%

E=3.2+(11.76-3.2)*1.62=17.0672%(or 17.07% approx).

A stock is correctly priced if required return=expected return

Hence Stock E is correctly priced.